# columns=colsep=4pt, rows=rowsep=.5pt

### v1

@@ C.1 ImageNet-1k
Method Backbone Dataset Epochs k-NN (%) Linear (%)
DINO (Caron et al., 2021) RN-50 ImageNet1kImageNet-1k 400 67.5 75.3
SWAV (Caron et al., 2020) RN-50 ImageNet1kImageNet-1k 400 65.0 74.6
CARP (Silva and Ramírez Rivera, 2023) RN-50 ImageNet1kImageNet-1k 400 67.7 75.3
CARP + Decoupling RN-50 ImageNet1kImageNet-1k 400 69.1 75.3
DINO (Caron et al., 2021) ViT-S/16 ImageNet1kImageNet-1k 300 72.8 76.2
iBOT (Zhou et al., 2022) ViT-S/16 ImageNet1kImageNet-1k 300 74.6 77.4
CARP + Decoupling ViT-S/16 ImageNet1kImageNet-1k 300 74.1 76.2
MSN (Assran et al., 2022) ViT-S/16 ImageNet1kImageNet-1k 600 76.9
iBOT (Zhou et al., 2022) ViT-S/16 ImageNet1kImageNet-1k 800 75.2 77.9
DINO (Caron et al., 2021) ViT-S/16 ImageNet1kImageNet-1k 800 74.5 77.0
DINO-vMF (Govindarajan et al., 2023) ViT-S/16 ImageNet1kImageNet-1k 800 74.7 77.0
iBOT-vMF (Govindarajan et al., 2024) ViT-S/16 ImageNet1kImageNet-1k 800 75.3 77.9
iBOT-vMF + KP (Govindarajan et al., 2024) ViT-S/16 ImageNet1kImageNet-1k 800 75.3 77.9
CARP + Decoupling ViT-S/16 ImageNet1kImageNet-1k 800 75.3 76.4
DINO (Caron et al., 2021) ViT-B/16 ImageNet1kImageNet-1k 400 76.1 78.2
DINO-vMF (Govindarajan et al., 2023) ViT-B/16 ImageNet1kImageNet-1k 400 77.4 78.8
iBOT (Zhou et al., 2022) ViT-B/16 ImageNet1kImageNet-1k 400 77.1 79.5
iBOT-vMF (Govindarajan et al., 2023) ViT-B/16 ImageNet1kImageNet-1k 400 78.7 80.3
iBOT-vMF + KP (Govindarajan et al., 2024) ViT-B/16 ImageNet1kImageNet-1k 400 78.8 80.5
CARP + Decoupling ViT-B/16 ImageNet1kImageNet-1k 400 76.7 78.1
C.2 iNaturalist2018
@@
DINO-vMF (Govindarajan et al., 2023) 1380 $(2.1\%)$ 49.7 68.5
iBOT-vMF (Govindarajan et al., 2023) 1804 $(22.0\%)$ 50.1 69.4
iBOT-vMF (kd) (Govindarajan et al., 2023; Oquab et al., 2024) 1843 $(22.5\%)$ 50.5 69.1
iBOT-vMF (kp) (Govindarajan et al., 2024) 7895 $(96.4\%)$ 51.1 69.3
MSN ($\lambda=1$) (Assran et al., 2022) 3363 $(41.3\%)$ 53.8 63.5
PMSN ($\lambda=5$) (Assran et al., 2023a) 3005 $(36.9\%)$ 41.8 64.2
@@
iBOT-vMF (kd) (Govindarajan et al., 2023; Oquab et al., 2024) 1634 $(19.9\%)$ 50.4 73.3
iBOT-vMF (kp) (Govindarajan et al., 2024) 7573 $(92.4\%)$ 51.4 74.0

### v2

@@ C.2.1 ImageNet-1k
Method Backbone Epochs k-NN (%) Linear (%)
DINO RN-50 $400$ 67.5 75.3
SWAV RN-50 $400$ 65.0 74.6
CARP RN-50 $400$ 67.7 75.3
CARP + Decoupling RN-50 $400$ 69.1 75.3
DINO ViT-S/16 $300$ 72.8 76.2
iBOT ViT-S/16 $300$ 74.6 77.4
CARP ViT-S/16 $300$ 73.6 76.3
CARP + KP ViT-S/16 $300$ 73.7 76.1
CARP + Decoupling ViT-S/16 $300$ 74.1 76.2
MSN ViT-S/16 $600$ 76.9
iBOT ViT-S/16 $800$ 75.2 77.9
DINO ViT-S/16 $800$ 74.5 77.0
DINO-vMF ViT-S/16 $800$ 74.7 77.0
iBOT-vMF ViT-S/16 $800$ 75.3 77.9
iBOT-vMF + KP ViT-S/16 $800$ 75.3 77.9
CARP + Decoupling ViT-S/16 $800$ 75.3 76.4
DINO ViT-B/16 $400$ 76.1 78.2
DINO-vMF ViT-B/16 $400$ 77.4 78.8
iBOT ViT-B/16 $400$ 77.1 79.5
iBOT-vMF ViT-B/16 $400$ 78.7 80.3
iBOT-vMF + KP ViT-B/16 $400$ 78.8 80.5
CARP + Decoupling ViT-B/16 $400$ 76.7 78.1
C.2.2 iNaturalist2018
@@
DINO-vMF 1380 $(2.1\%)$ 49.7 68.5
iBOT-vMF 1804 $(22.0\%)$ 50.1 69.4
iBOT-vMF (kd) 1843 $(22.5\%)$ 50.5 69.1
iBOT-vMF (kp) 7895 $(96.4\%)$ 51.1 69.3
MSN ($\lambda=1$) 3363 $(41.3\%)$ 53.8 63.5
PMSN ($\lambda=5$) 3005 $(36.9\%)$ 41.8 64.2
@@
iBOT-vMF (kd) 1634 $(19.9\%)$ 50.4 73.3
iBOT-vMF (kp) 7573 $(92.4\%)$ 51.4 74.0
@@
C.3 Sensitivity of DINO with full decoupling
Table C.3: Sensitivity analysis of DINO under different configurations of hyperparameters.
Method Dec. Cent. SK # Prototypes Teacher temperature k-NN (%)
DINO $\checkmark$ 65536 $0.04\rightarrow 0.07$ 68.9 $\downarrow$  0.
[…]
DINO $\checkmark$ $\checkmark$ 65536 $0.04\rightarrow 0.07$ $2\downarrow  6.$71
[…]
Our re-produced DINO backbone using 6 local crops and mixed precision training.
While the proposed decoupling method mitigates prototypical collapse in both CARP (Silva and Ramírez Rivera, 2023) and DINO (Caron et al., 2021), it also introduces training instabilities in DINO. To examine DINO’s sensitivity to increased prototype diversity, we reproduce a baseline DINO model using the official codebase with two slight modifications. First, we adapt the dataset loader to meet the constraints of our High-Performance Computing (HPC) environment, which prevents us from reusing the exact same data-loading seed. Second, to reduce training time and computational cost, we employ mixed-precision training instead of float32.
The resulting k-NN performance is reported in Table C.3. While the original DINO paper reports a k-NN accuracy of 69.7, our reproduced model exhibits a modest performance drop of 0.8 percentage points. We attribute this difference primarily to (a) the altered data ordering introduced by our changed dataloader and (b) the reduced numerical precision during training. All other baselines are produced under the same conditions, ensuring that the results in the table remain directly comparable. All models are trained on ImageNet-1k for 100 epochs using six local crops.
Building on these observations, we next investigate how DINO behaves when its prototype optimization is replaced by the decoupled online GMM update. As shown in Table C.3, replacing the prototype update mechanism from gradient descent to an online GMM, while keeping all other hyperparameters fixed, results in a performance reduction of 41.1 percentage points. This suggests that the centering operation in the standard DINO formulation is insufficient to promote sufficiently diverse assignments when prototype diversity increases. To address this limitation, we employ Sinkhorn-Knopp normalization (Cuturi, 2013), which substantially improves assignment uniformity and reduces the performance gap with vanilla DINO to 6.7 percentage points. This effect on improving uniformity accross assignments was first identified by Ruan et al. (2023) and later adopted by Oquab et al. (2024) in the DINOv2 framework.
Next, we follow the recommendations of Ruan et al. (2023) and adopt their teacher-temperature scheduling strategy, which progressively increases the sharpening of the teacher’s assignment distribution and thus provides a stronger learning signal. We observe that this modification substantially benefits DINO under increased prototype diversity: the additional sharpening yields a further improvement of 5.5 percentage points, narrowing the remaining gap to the vanilla DINO baseline even more effectively.
Finally, we investigate whether adjusting the number of prototypes can further stabilize training under the decoupled formulation. We find that substantially reducing the prototype count to 1k yields a small additional improvement, leaving a remaining gap of 0.9 percentage points to the baseline. Exploring this further, we identify an intermediate setting with 2k prototypes as the most effective configuration: this choice closes the gap almost entirely, reducing the difference to only 0.2 percentage points.
We use these selected configurations for all DINO + Decoupling experiments in Section 4.3 and Section 4.2. For the long-tailed setting in Section 4.4, we retain the same hyperparameters but increase the number of prototypes to 8192 to match the 8142 classes in iNaturalist-18. Applying the settings tuned on ImageNet-1k to iNaturalist-18 leads to a substantial performance drop, as reported in Table 2. Although further tuning tailored to the long-tailed distribution could likely mitigate this behavior, such investigation falls outside the scope of this work.
Ultimately, we find the pronounced sensitivity of DINO to prototype diversity to be an interesting empirical observation. Although CARP is architecturally similar to DINO, it differs in four key components: random partitioning to stabilize training, unnormalized projection-head outputs, entropy regularization in place of centering the teacher logits, and a loss formulation that favors more one-hot-like predictions. Any one of these design choices, or some combination thereof, may underlie CARP’s ability to remain stable and benefit from increased prototype diversity. Determining which components are responsible for this robustness, and how they interact, represents a compelling direction for future research.
Appendix D Declaration of Use of Large Language Models
During the preparation of this work, the author(s) used Large Language Models (LLMs) for grammar and spelling checks, paraphrasing and rewording, code writing, writing assistance and reviewing, as well as for identifying and summarizing related work. After using these services, the author(s) reviewed and edited all generated content as needed and take(s) full responsibility for the publication’s content.